# CLIMB-3D: Continual Learning for Imbalanced 3D Instance Segmentation

### v2

@@ 3.3 Benchmarking Incremental Scenarios
While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution – an assumption that rarely holds in real-world settings. In practice, the number of object categories, denoted by $\mathcal{C}$, is often large and characterised by substantial variability in category frequency, shape, structure, and size. To address these challenges, we propose three incremental learning scenarios, each designed to capture a different facet of real-world complexity. The design of these scenarios is detailed below; for more information and illustration, please refer Appendix Appendix C of the supplementary material.
Split_A: Frequency Scenarios. This scenario acknowledges that datasets are often labelled based on the frequency of category occurrences. To accommodate this, we propose a split where the model initially learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritizing the training of frequently occurring categories, the model can establish a strong foundation before expanding its knowledge to handle rare categories.
Split_B: Semantic Scenarios. Beyond frequency, semantic similarity is crucial in real-world deployments. While objects often share visual or functional traits, models may encounter semantically different categories in new environment. To simulate this challenge, we introduce the Split_Bscenario. Here, categories are grouped based on their semantic relationships, and the model is incrementally trained on one semantic group at a time. This setup encourages the model to generalise across semantically similar categories and adapt more effectively when exposed to new ones. Unlike the Split_Ascenario, which organises learning based on category frequency, the Split_Bscenario may contain both frequent and infrequent categories within a single task, focusing instead on semantic continuity and transfer.
Split_C: Random Scenarios. In some cases, data labelling is driven by the availability of objects rather than predefined criteria. To capture this, we introduce the Split_Cscenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance. By exposing the model to such diverse and unpredictable distributions, we aim to improve its robustness in real-world situations where labelled data availability is inconsistent.
By designing these incremental scenarios, we aim to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
@@ 4 Experiments
Split_A CLIMB-3D (Ours) 35.69 31.05 22.72 3.44 2.63
@@
Split_B CLIMB-3D (Ours) 35.48 31.56 23.69 8.00 5.51
@@
Split_C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
@@
Table 3: Ablation study showing the impact of each component ER, PLG, and CBR in a three-phase setup. Each split (s) and number corresponds to data introduced at that phase. The All column reports performance over all classes seen so far. Joint training (Oracle) is highlighted in gray, and the full model in orange. Best results are in bold.
@@
4.1 Setup
Datasets. We evaluate our method on the ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance – making it ideal for simulating and assessing real-world scenarios. In addition, we benchmark our approach against existing incremental learning methods using the ScanNetV2 [dai2017scannet] in a 3DSS setting. For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR].
Evaluation Metrics. We evaluate our method using mean Average Precision (mAP), a standard metric for 3DIS that provides a comprehensive measure of segmentation quality by accounting for both precision and recall. For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments. To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual]. FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
Incremental Scenarios. As discussed in Section 3.3, we design three incremental scenarios – Split_A, Split_B, and Split_C– each with three tasks, based on object frequency, semantic similarity, and random grouping, respectively. For Split_A, we follow the class split provided by rozenberszki2022language, with a class distribution of 66-68-66 across the three tasks. In Split_B, we partition the classes into 74-50-76 based on semantic similarity, which is calculated using the CLIP [radford2021learning] text encoder, followed by clustering using K-Means. Finally, in Split_C, classes are randomly shuffled and split into three sets, with 67-67-66 categories per split. These scenarios comprehensively evaluate our approach’s performance and generalization in varied real-world conditions.
Implementation Details. For all the experiments we use Mask3D [Schult23ICRA] with the Minkowski Res16UNet34C [choy20194d] as a feature backbone. We adopt the data augmentation, hyperparameters, and training strategy described in [Schult23ICRA], with using the AdamW optimizer [loshchilov2017decoupled] and a one-cycle learning rate scheduler [smith2018superconvergence], and results are evaluated on the entire validation set. In incremental training, learning is divided into three phases, introducing one split per phase across the three scenarios described in Sec. 3.3. For each phase, we train for 500, 300, and 300 epochs, respectively, except for the Oracle setting (Row 1 in Tab. 3), which is trained once for 600 epochs. We retain the same hyperparameters throughout, modifying only the number of training epochs. The exemplar replay size is fixed at $|\mathcal{E}^{t}|=50$ scenes.
4.2 Results and Discussion
To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task. In contrast, for the 3DSS setting, where prior incremental baselines do exist, we compare our method against three existing approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
Results on CI-3DIS. Tab. 1 compares CLIMB-3D against the ER baseline across three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP. In the Split_Ascenario, characterized by significant distribution shifts, it achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, while drastically reducing forgetting with FPP scores of 3.44% and 2.63%, over 47 points lower than the baseline. Under the Split_Bscenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%. For the Split_Cscenario, featuring increasing geometric complexity, it maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points. Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, validating the method’s effectiveness for scalable continual 3D semantic understanding.
Results on CI-3DSS. Although our primary focus is on CI-3DIS, we additionally evaluate our method under the CI-3DSS setting to enable comparison with existing approaches. For this, we adapt our predictions by assigning each point the label corresponding to the highest-confidence mask and exclude background classes (floor and wall), as these are not part of the semantic segmentation targets. Tab. 2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall. Despite being designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains: +35.23% mIoU in Phase 1 and +19.1% in Phase 2. Overall, it achieves 59.38% mIoU – significantly higher than the $\sim$30mIoU of prior methods – highlighting its robustness and transferability.
For a detailed per-phase analysis, evaluation of rare classes, and qualitative result comparisons, please refer to Appendix Appendix A, Appendix Appendix B, and Appendix Appendix D, respectively, in the supplementary material.
4.3 Ablation
@@
We conduct an ablation study to evaluate the contribution of each component in our framework. As an upper bound, we report the performance of an Oracle model trained jointly on the full dataset. For the incremental setting, we follow the previously defined splits and first train the model naively across phases. We then incrementally add each module to isolate its impact. Tab. 3 presents results on the Split_Ascenario using both $\text{mAP}_{50}$  and FPP metrics.
@@
Naïve Training. When trained naïvely without any dedicated modules, the model suffers from severe catastrophic forgetting, as evident in row 2. The model completely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and a high FPP of 56.82.
@@
Effect of PLG. The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from earlier phase, facilitates knowledge retention and forward transfer. As shown in row 4, PLG significantly reduces forgetting and enhances performance on current tasks. For s1, it improves $\text{mAP}_{50}$ by 31.49% in phase 2 and 39.40% in phase 3 over exemplar replay. Overall, PLG yields a 15.00% increase in performance and reduces forgetting by 39.40%.
Effect of CBR. Finally, class-balanced re-weighting (CBR) mitigates class imbalance both during pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency. As shown in row 5, CBR enhances s1 retention over PLG in phases 2 and 3 (+4.67% and +4.41%) and further improves performance on s2 and s3. It also achieves the lowest FPP of 2.63, reflecting effective forgetting mitigation. Overall, CBR offers the best trade-off between retaining past knowledge and learning new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
@@ 5 Conclusion
We address the critical challenge of catastrophic forgetting in class-incremental 3D instance segmentation by proposing a modular framework that combines exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR). Our method incrementally adapts to new classes while preserving knowledge from previous tasks without requiring access to the full dataset. Through extensive experiments on a three-phase benchmark, we demonstrate the individual and combined effectiveness of each component. ER provides a easy knowledge storage, PLG enhances knowledge retention via pseudo-label supervision, and CBR further refines learning by mitigating class imbalance. Together, these modules significantly reduce forgetting and improve overall segmentation performance across all phases. By balancing stability and plasticity, our method contributes to continual 3D scene understanding and establishing a new baseline for future research.
@@
Table 4: Comparison of results in terms of mAP50 with and proposed CLIMB-3D for three different scenarios. Each scenario is trained in three phases ($\mathtt{phase}=1,2,3$) by introducing a single split s at a time. The results highlighted in orange are with the proposed method, and the best results for each scenario are in bold.
@@
Split_C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split_A  for Phase 2, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER+KD ER+KD+IC
@@
Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split_A  for Phase 3, in terms of $\text{mAP}_{50}$.
@@
Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels. The three scenarios Split_A, Split_A, and Split_C  are each divided into three tasks: Task 1, Task 2, and Task 3.
@@
In this supplementary material, we first present the per-phase performance of the proposed CLIMB-3D model in Appendix Appendix A, followed by the analysis of gains on rare classes achieved through the incorporation of the CBR module in Appendix Appendix B. Next, we provide detailed split information for all scenarios, based on class names, in Appendix Appendix C. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix Appendix D.
Appendix Appendix A Per-Phase Performance Analysis
We extend the analysis presented in Tab. 1 (Sec. 4.2 of the main paper) to Tab. 4 to examine the impact of our proposed method across individual splits and phases in different scenarios. The results consistently show that our method outperforms the baseline by better retaining knowledge from earlier tasks.
Split_A. In the Split_Ascenario, our model shows notable improvements across all phases. Particularly in Phase 3, although both methods show a drop in performance for s2, our approach significantly reduces forgetting, maintaining a higher $\text{mAP}_{50}$  across all splits.
Split_B. The Split_Bscenario, while more challenging due to increased semantic overlap, demonstrates our model’s strength in generalisation. In Phase 2, it achieves an overall $\text{mAP}_{50}$  of 43.13% compared to the baseline’s 24.53%. This trend continues in Phase 3, with our model achieving 31.56%, more than double the baseline’s 15.07%, indicating both improved learning and retention.
Split_C. For the Split_Cscenario, where splits are randomly grouped, Phase 1 performance is initially lower. However, in subsequent phases, our method balances stability and plasticity better than the baseline. By Phase 3, it consolidates performance across all splits, improving the overall $\text{mAP}_{50}$to 26.78%, a 5.7% gain over the baseline.
These results demonstrate the robustness of our method in handling continual learning under varied and challenging split strategies.
Appendix Appendix B Evaluation on Rare Categories
The proposed Class-Balanced Re-weighting (CBR) module, as detailed in 3.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and Pseudo-Label Generator (PLG). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
The results, shown in Tab. 5 and Table 6, correspond to evaluations on Split_A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Table 5, the CBR module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split_A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, show significant improvement when the CBR module is applied. Overall, the CBR module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
Similarly, Table 6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though classes such as piano, bucket, and laundry basket are observed only once, the CBR module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+PLG module does not focus on rare classes like shower and toaster, which results in low performance, but the CBR module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed CBR module into the framework outperforms ER+PLG by 12.13%.
Appendix Appendix C Incremental Scenarios Phases
@@
Table 7 presents the task splits for each proposed scenario introduced in Sec. 3.3, and Fig. 3 illustrates them visually for the ScanNet200 [rozenberszki2022language] dataset. The three scenarios – Split_A, Split_B, and Split_C– are each divided into three tasks: Task1, Task2, and Task3. Notably, the order of classes within these tasks is random.
Appendix Appendix D Qualitative Results
@@
In this section, we present a qualitative comparison of the proposed framework with the baseline method. Figure 4 illustrates the results on the Split_A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach. As shown in the figure, our method demonstrates superior instance segmentation performance compared to the baseline. For example, in row 1, the baseline method fails to segment the sink, while in row 3, the sofa instance is missed. Overall, our framework consistently outperforms the baseline, with several missed instances by the baseline highlighted in red circles.
@@
Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_A  evaluation after learning all tasks.
@@
In Figure 5, we present the results on Split_B, highlighting instances where the baseline method underperforms, marked with red circles. For example, in row 2, the baseline method incorrectly identifies the same sofa as separate instances. Similarly, in row 5, the washing machine is segmented into two instances by the baseline. In contrast, the proposed method delivers results that closely align with the ground truth, demonstrating its superior performance
@@
Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_B  evaluation after learning all tasks.
@@
Similarly, Figure 6 highlights the results on Split_C, where classes are encountered in random order. The comparison emphasizes the advantages of our method, as highlighted by red circles. The baseline method often misses instances or splits a single instance into multiple parts. In contrast, our approach consistently produces results that are closely aligned with the ground truth, further underscoring its effectiveness.
@@
Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split_C  evaluation after learning all tasks.

### v3

@@ 3.3 Benchmarking Incremental Scenarios
While CI-3DIS methods offer a wide range of practical applications, they frequently rely on the assumption of uniform sample distribution, which rarely holds in real-world settings. In practice, the number of object categories, denoted by $\mathcal{C}$, is often large and characterised by substantial variability in category frequency, shape, structure, and size. To address these challenges, we propose three incremental learning scenarios, each designed to capture a different facet of real-world complexity. The design of these scenarios is detailed below; for further information and illustrations, refer to Appendix E of the supplementary material.
Split-A: Frequency Scenarios. This scenario acknowledges that datasets are often labelled based on the frequency of categories. To accommodate this, we propose a split where the model learns from the most frequent categories and subsequently incorporates the less frequent ones in later stages. By prioritising the training of frequently occurring categories, the model can establish a strong foundation before expanding to handle rare categories.
Split-B: Semantic Scenarios. Beyond frequency, semantic similarity is crucial in real-world deployments. While objects often share visual or functional traits, models may encounter semantically different categories in new environments. To simulate this challenge, we introduce the Split-B scenario. Here, categories are grouped based on their semantic relationships, and the model is incrementally trained on one semantic group at a time. This setup encourages the model to generalise across semantically similar categories and adapt more effectively when exposed to new ones. Unlike the Split-A scenario, which organises learning based on category frequency, the Split-B scenario may contain both frequent and infrequent categories within a single task, focusing on semantic continuity and transfer.
Split-C: Random Scenarios. In some cases, data labelling is driven by the availability of objects rather than predefined criteria. To capture this, we introduce the Split-C scenario, which represents a fully random setting where any class can appear in any task, resulting in varying levels of class imbalance. By exposing the model to such diverse and unpredictable distributions, we aim to improve its robustness in real-world situations where labelled data availability is inconsistent.
These incremental scenarios are designed to provide a more realistic representation of object distributions, frequencies, and dynamics encountered in the real world.
@@ 4 Experiments
4.1 Setup
Datasets. We evaluate CLIMB-3D on ScanNet200 [rozenberszki2022language], which comprises 200 object categories and exhibits significant class imbalance, which makes it ideal for simulating and assessing real-world scenarios. In addition, we compare our approach against existing incremental learning methods using ScanNetV2 [dai2017scannet] in the 3DSS setting. For this evaluation, we follow the standard training and validation splits defined in prior works [Yang_2023_CVPR] to ensure consistency with existing methods.
Evaluation Metrics. We evaluate our method using Mean Average Precision (mAP), a standard metric for 3DIS which provides a comprehensive measure of segmentation quality by accounting for both precision and recall. For comparison with existing 3DSS methods, we report the mean Intersection over Union (mIoU), which quantifies the overlap between predicted and ground truth segments. To assess the model’s ability to mitigate catastrophic forgetting in incremental settings, we use the Forgetting Percentage Points (FPP) metric, as defined in [liu2023continual]. FPP captures performance degradation by measuring the accuracy drop on the initially seen categories between the first and final training phases.
Detailed descriptions of the incremental scenarios (Split-A, Split-B, Split-C) and implementation are provided in Appendix B.
4.2 Results and Discussion
@@
Split-A CLIMB-3D (Ours) 35.69 31.05 22.72 3.44 2.63
@@
Split-B CLIMB-3D (Ours) 35.48 31.56 23.69 8.00 5.51
@@
Split-C CLIMB-3D (Ours) 31.59 26.78 18.93 9.10 7.89
To evaluate our proposed approach, we construct a baseline using ER [buzzega2020dark] for the 3DIS setting, as no existing incremental baselines are available for this task. In contrast, for the 3DSS setting, where incremental baselines do exist, we compare our method against three prior approaches [serra2018overcoming, Yang_2023_CVPR, li2017learning].
Results on CI-3DIS. Tab.˜1 compares CLIMB-3D against the ER baseline across the three CI-3DIS scenarios, evaluated after all phases using $\text{mAP}_{25}$, $\text{mAP}_{50}$, overall mAP, and FPP. In the Split-A scenario, characterised by significant distribution shifts, our method achieves gains of +19.23%, +16.76%, and +12.28% in $\text{mAP}_{25}$, $\text{mAP}_{50}$, and overall mAP, respectively, while drastically reducing forgetting with FPP scores of 3.44% ($\text{mAP}_{25}$) and 2.63% ($\text{mAP}_{50}$), both over 45 points lower than the baseline. Under the Split-B scenario, with semantically related new categories, improvements of +18.26%, +16.49%, and +12.76% are observed alongside a reduction in forgetting to 8.00% and 5.51% from baseline levels above 42%. For the Split-C scenario, featuring increasing geometric complexity, CLIMB-3D maintains solid gains of +5.94%, +5.70%, and +4.08% across mAP metrics, with forgetting reduced by over 22 points. Overall, these results demonstrate robust forward transfer, minimal forgetting, and stable learning, highlighting the method’s effectiveness for scalable continual 3D semantic understanding.
Results on CI-3DSS.
@@
Although our primary focus is on CI-3DIS, we also evaluate our method under the CI-3DSS setting to enable comparisons with existing approaches. To do so, we adapt our predictions to assign each point the label corresponding to the highest-confidence mask and exclude background classes (e.g., floor and wall), as these are not part of the semantic segmentation targets. Tab.˜2 presents a comparison between our method and existing baselines on the ScanNet V2 dataset [dai2017scannet], evaluated using mIoU across two training phases and overall. Although originally designed for instance segmentation, our method generalises effectively to semantic segmentation, achieving substantial gains of +35.23% mIoU in Phase 1 and +19.1% in Phase 2. Overall, it achieves 59.38% mIoU, significantly higher than the $\sim$30% mIoU of prior methods, highlighting its robustness and transferability.
Refer to Appendix C, D, and F of the supplementary material for detailed analyses of per-phase performance, rare-class evaluation, and qualitative result comparisons, respectively.
4.3 Ablation
Table 3: Ablation study illustrating the impact of each component in a three-phase setup. Each split (s) and corresponding number indicates data introduced at that phase. Best results are highlighted in bold.
@@ 4.3 Ablation
We conduct an ablation study to evaluate the contribution of each component in our framework. As an upper bound, we report the performance of an Oracle model, which is trained jointly on the full dataset. For the incremental setting, we follow the previously defined splits and first train the model naively across phases. We then incrementally add each module to isolate its impact. Tab.˜3 presents results for the Split-A scenario using both $\text{mAP}_{50}$  and FPP metrics.
@@
Naïve Training. When trained naïvely without any dedicated modules, the model suffers from severe catastrophic forgetting, as evident in row 2. The model entirely forgets previously learned classes upon entering new phases, with performance dropping to zero for earlier splits and FPP rising to 56.82.
@@
Effect of PLG. The addition of the pseudo-label generator (PLG), which generates labels for previous classes by retaining a copy of the model from an earlier phase, facilitates knowledge retention and forward transfer. As shown in row 4, PLG significantly reduces forgetting and enhances performance on current tasks. For s1, it improves $\text{mAP}_{50}$ by 31.49% in Phase 2 and 39.40% in Phase 3 over exemplar replay. Overall, PLG yields a 15.00% increase in performance and reduces forgetting by 39.40%.
Effect of CBR. Finally, class-balanced re-weighting (CBR) mitigates class imbalance during both pseudo-labelling and current task learning by adjusting each class’s contribution based on its frequency. As shown in row 5, CBR enhances s1 retention over PLG in Phases 2 and 3 (+4.67% and +4.41%), and further improves performance on s2 and s3. It also achieves the lowest FPP of 2.63, reflecting strong forgetting mitigation. Overall, CBR offers the best trade-off between retaining past knowledge and acquiring new tasks, outperforming PLG and ER by 4.41% and 43.81%, respectively.
4.4 Limitations & Future Work
While CLIMB-3D achieves strong performance across CI-3DIS and CI-3DSS settings, several limitations remain and warrant discussion. The experiments are currently limited to indoor datasets (ScanNet200 and ScanNetV2), therefore evaluating performance on outdoor scenes would help assess generalisability across diverse environments. The current setup also considers only three incremental phases, whereas real-world applications often involve longer sequences; incorporating such extended setups would better reflect practical challenges. Moreover, the model operates in a uni-modal setting. Integrating multi-modal cues, such as vision-language models, could further enhance performance, as suggested by recent 2D and 3D studies [hong20233d, thengane2022clip, thengane2025foundational].
@@ 5 Conclusion
This work addresses the challenge of catastrophic forgetting in class-incremental 3D instance segmentation by introducing a modular framework that integrates exemplar replay (ER), a pseudo-label generator (PLG), and class-balanced re-weighting (CBR). The proposed method incrementally adapts to new classes while preserving prior knowledge, all without requiring access to the full dataset. Extensive experiments on a three-phase benchmark validate the individual and combined effectiveness of each component: ER enables efficient memory-based retention; PLG facilitates knowledge preservation through pseudo-supervision; and CBR mitigates class imbalance to enhance learning stability. Together, these components significantly reduce forgetting and improve segmentation performance across all phases. By achieving a strong balance between stability and plasticity, our approach advances continual 3D scene understanding and sets a new baseline for future research in this area.
@@
Appendix Appendix A Illustrating Incremental scenarios
Appendix Appendix B Per Phase Analysis
Table 4: Comparison of results in terms of mAP50 with the proposed CLIMB-3D across three different scenarios. Each scenario is trained over three phases ($\mathtt{phase}=1,2,3$), introducing a single split s at a time. Results highlighted in orange correspond to the proposed method, and the best results for each scenario are shown in bold.
@@
Split-C CLIMB-3D 36.40 32.63 33.38 33.00 28.51 17.11 34.64 26.78
We extend the analysis from Tab.˜1 to Tab.˜4 to highlight the impact of our proposed method on individual splits across various scenarios. The results clearly demonstrate that our model consistently retains knowledge of previous tasks better than the baseline. For Split-A, our model shows improvement throughout the phase. In Phase 3 of (s2), although both the baseline and our method exhibit a performance drop, our method reduces forgetting significantly compared to the baseline. The Split-B scenario, while more complex than Split-A, achieves comparable results due to semantic similarity among classes within the same task. In Phase 2, our model achieves overall all 43.13% $\text{mAP}_{50}$ compared to 24.53% on baseline, a similar trend is observed in Phase 3, where our method not only consistently improves learning but also enhances retention of previous information. After all three tasks, our method achieves an overall performance of 31.56% AP50, compared to 15.07% for the baseline. In the Split-C scenario, the first-stage model struggles due to the increased complexity introduced by random grouping. In Phase 2, while the baseline focuses on learning the current task, it suffers from severe forgetting of prior knowledge. Conversely, our method balances new task learning with the retention of earlier information. By Phase 3, the model effectively consolidates s1 and maintains strong performance across all task splits. Overall, our proposed method improves mAP by 5.6%.
In this supplementary material, we first demonstrate the performance gains on rare classes achieved by incorporating the IC module in Appendix C. Next, we provide detailed split information for all scenarios, based on class names, in Appendix D. Finally, we present a qualitative comparison between the baseline method and our proposed approach in Appendix E.
Appendix Appendix C Evaluation on Rare Categories
The proposed imbalance correction (IC) module, as detailed in Section 4.2, is designed to address the performance gap for rare classes. To assess its impact, we compare its performance with the framework which has exemplar replay (ER) and knowledge distillation (KD). Specifically, we focus on its ability to improve performance for rare classes, which the model encounters infrequently compared to more common classes.
Table 5: Results for classes observed by the model 1–20 times during an epoch, evaluated on Split-A  for Phase 2, in terms of $\text{mAP}_{50}$.
Classes Seen Count ER + PLG ER + PLG + CBR
@@
The results, shown in Tab.˜5 and Tab.˜6, correspond to evaluations on Split-A  for Phase 2 and Phase 3, respectively. In Phase 2, we evaluate classes seen 1–20 times per epoch, while Phase 3 targets even less frequent classes, with observations limited to 1–10 times per epoch.
As illustrated in Tab.˜5, the IC module substantially improves performance on rare classes in terms of $\text{mAP}_{50}$  in Phase 2 of Split-A. For instance, classes like recycling bin and trash bin, seen only 3 and 7 times, respectively, shows significant improvement when the IC module is applied. Overall, the IC module provides an average boost of 8.32%, highlighting its effectiveness in mitigating class imbalance.
Table 6: Results for classes observed by the model 1–10 times during an epoch, evaluated on Split-A  for Phase 3, in terms of $\text{mAP}_{50}$.
@@
Similarly, Tab.˜6 presents results for Phase 3, demonstrating significant gains for infrequent classes. For example, even though the classes such as piano, bucket, and laundry basket are observed only once, IC module improves the performance by 52.30%, 10.40%, and 13.60%, respectively. The ER+KD module does not focus on rare classes like shower and toaster which results in low performance, but the IC module compensates for this imbalance by focusing on underrepresented categories. On average, the addition of the proposed IC module into the framework outperforms ER+KD by 12.13%.
Table 7: Classes grouped by tasks for each proposed scenario on the ScanNet200 dataset labels. The three scenarios Split-A, Split-A, and Split-C  are each divided into three tasks: Task 1, Task 2, and Task 3.
@@
Appendix Appendix D Incremental Scenarios Phases
@@
Tab.˜7 presents the task splits for each proposed scenario introduced in Section 4.3 using the ScanNet200 dataset. The three scenarios, Split-A, Split-B, and Split-C, are each divided into three tasks: Task 1, Task 2, and Task 3. Notably, the order of classes in these tasks is random.
Appendix Appendix E Qualitative Results
@@
In this section, we present a qualitative comparison of the proposed framework with the baseline method. Fig.˜4 illustrates the results on the Split-A evaluation after learning all tasks, comparing the performance of the baseline method and our proposed approach. As shown in the figure, our method demonstrates superior instance segmentation performance compared to the baseline. For example, in row 1, the baseline method fails to segment the sink, while in row 3, the sofa instance is missed. Overall, our framework consistently outperforms the baseline, with several missed instances by the baseline highlighted in red circles.
@@
Figure 4: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split-A  evaluation after learning all tasks.
@@
In Fig.˜5, we present the results on Split-B, highlighting instances where the baseline method underperforms, marked with red circles. For example, in row 2, the baseline method incorrectly identifies the same sofa as separate instances. Similarly, in row 5, the washing machine is segmented into two instances by the baseline. In contrast, the proposed method delivers results that closely align with the ground truth, demonstrating its superior performance
@@
Figure 5: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split-B  evaluation after learning all tasks.
@@
Similarly, Fig.˜6 highlights the results on Split-C, where classes are encountered in random order. The comparison emphasizes the advantages of our method, as highlighted by red circles. The baseline method often misses instances or splits a single instance into multiple parts. In contrast, our approach consistently produces results that are closely aligned with the ground truth, further underscoring its effectiveness.
@@
Figure 6: Qualitative comparison of ground truth, the baseline method, and our proposed framework on the Split-C  evaluation after learning all tasks.